Value of a for which the given equation
logx/loga = a^a^x has one solution..

Grade:11

2 Answers

bharat bajaj IIT Delhi
9 years ago
Log x/log a= a^a^x
x = a^a^a^x

There is one possible solution when a= 1
Hence x = 1^x
x=1 is the solution...

Thanks
Bharat Bajaj
Askiitians faculty
Arun Kumar
9 years ago
Hi bharat bajaj 
for a=1 LHS would’t be defined because log(a)=0 !!!!!